# Jointly modelling the evolution of community structure and language in online extremist groups

### v1

@@ 4 Data
As a background corpus to calculate the PPMI of the content matrices, we use a dataset of over 40 million Reddit posts by Dziri et al. [11].
5 Experimental setup
We run preliminary experiments on a reduced training set to establish the efficacy of the optimisations outlined in Section 3.4 and to perform hyperparameter tuning. These results, reported in Appendix A, show that the optimisations provide statistically significant improvements in the reconstruction accuracy in a smaller scale evaluation setting, and as such they are incorporated in all models that are compared in this work.
Following Appel et al. [2], we use two test cases for evaluation: clustering and forecasting.
@@ 5.1 Clustering evaluation
We use K-means clustering based on the temporal user embeddings $U$ to discover sub-communities. Following Appel et al. [2], these clusters are calculated across different timesteps, meaning that a user may belong to the same cluster or to different clusters over different timesteps. Intuitively, the clusters may be thought of as topics, such that users who do not overlap temporally but have similar embeddings at different timesteps may be assigned to the same cluster.
Using the subreddit structure, we evaluate the soundness of these clusters based on cluster purity, which evaluates uniformity of cluster elements as the fraction of users in each cluster that belong to the dominant class [22]. Since one user may be associated with multiple subreddits at a given timestep, we use a multilabel interpretation of this metric, meaning that the dominant class is the class that has the highest representation in a cluster. Let $L(i,t)$ denote the label set of user $i$ at time $t$, $\Omega=\{\omega_{1},\omega_{2},\ldots,\omega_{K}\}$ the set of clusters, and =$\mathbb{C}=\{c_{1},c_{2},\ldots,c_{J}\}$ the set of classes. We report the average cluster purity over $K$ clusters, given:
@@
The class labels comprise a two-level hierarchy, consisting of a more general category label (with 7 classes: Incels, MGTOW, TRP, PUA, MRA, mental health and criticism) and a more finegrained subreddit label (with 50 classes). Allowing for some subgroup specialisation, we report results for $k\in\{10,100,1000\}$, using both subreddit and category-level labels. Since the clustering algorithm is dependent on random initialisation, we report the mean and standard deviation over 5 runs.
@@ 5.3 Community prediction
We use a distance-based approach for this purpose. We find the centroid of each community, i.e. the mean of the embeddings of all users who interacted with said community in $t_{1}$ to $t_{T}$. We then calculate the cosine similarity $s$ between the predicted user embedding and the centroid of each community. The inferred user label is given by the max-normalised similarities to all centroids, ie. $\hat{y_{i,t}}=\frac{s_{i,t}}{max(s_{i,t})}$, to capture the relative relevance of each community to a user.
@@
This formulation equates to a multilabel regression task, which we evaluate using the concordance index (CI) [17]. The CI is the fraction of concordant predictions (ie.  ($y_{i}>y_{j}$, $\hat{y}_{i}>\hat{y}_{j}$) or ($y_{i}<y_{j}$, $\hat{y}_{i}<\hat{y}_{j}$); in other words, a pairwise comparison is performed on all possible pair permutations, and the fraction that are correctly ordered is calculated. We report both the per-class CI (comparing predictions across users for a specific community) and the within-sample CI (evaluating the orderings per user).
@@ 6.1 Cluster purity
K=10 K=100 K=1000
Model Category Subreddit Category Subreddit Category Subreddit
Cluster purity over $K$ clusters is shown in Table 3. Our model outperforms the StatCont and NoAdj systems in 4 out of 6 cases, with the largest improvement observed for subreddit purity at $K=1000$. The NoAdj model has the same score as ours for category purity at $K=100$, whereas the StatCont model performs the same as ours for category purity at $K=1000$. Our model has a larger standard deviation for 5/6 cases, which may be due to an increased flexibility through a larger number of parameters.
We note that cluster purity is better at larger $K$, which is to be expected; purity will be 1 if $K=N$. In all cases, the cluster purity is higher for the category labels than the subreddit labels, which again is to be expected; since users are distributed over more classes, clusters are less likely be dominated by a single subreddit.
We also compare our cluster purity to embeddings that are obtained through standard static (i.e. time-aggregated) matrix factorisation of the content embedding (MatFact) and static shared matrix factorisation of content and adjacency (SharedMF). In this case, only one embedding is produced per user and therefore a user is assigned to only one cluster. We calculate purity based on all clusters with which a user interacted over the training period. From the results, we note that the cluster purity is lower in this setting. A possible reason for this result is that users vary in their category associations over time, which is not captured well by the time-aggregated clustering. The static shared matrix factorisation model outperforms the static model without social information.
The per-cluster purity for $K=10$, at the category level and using Cerberus, is shown in Table 2, sorted by cluster size. Size is based on the number of (user, timestep) tuples in the cluster. We note that there is a substantial difference between the high-scoring and low-scoring clusters, with clusters 2, 7, 8, 9 and 10 being high-purity clusters representing 4 out of the 6 higher level groups in the data. With the exception of cluster 2, these are all relatively small clusters, possibly representing the dedicated core users of each ideology who do not engage with other communities. Cluster 2 is a large and high-purity cluster which represents the mental health category. This provides support for the soundness of the clustering and embeddings, as we would expect this control group to be disparate.
@@ 6.2 Evolution of users
Figure 2: Predicting user evolution over time. Grey dots represent the position of a given user, whereas coloured dots are community centroids. At each timestep, the predicted next embedding is shown in red and the true next embedding in black.
In the context of extremist groups, the ability to represent a user as a dynamic entity at different timesteps is a key benefit of the temporal factorisation approach. Prior work [9, 14] developed predictive models to anticipate whether a user will engage with an extremist group; however, these works relied on derived indicators, whereas we our system models users and communities in the same embedding space.
@@
As described in Section 5, we evaluate user evolution modelling in two ways: embedding prediction (comparing learned embeddings to predicted embeddings, within the training window) and community prediction (predicting beyond the training window and comparing to true interactions). These results of are shown in Table 4. Though we use the subreddit structure to evaluate our models, they can be applied in forums without explicit subcommunities.
@@ 7 Evolution of words
Figure 3: Word relevance to different clusters over time. Labels are assigned based on the majority class in the cluster. Linewidth represents the cluster size.
@@
There are three main takeaways from these figures. Firstly, dynamic word embeddings are important: a static model would produce an aggregate representation of Kavanaugh, which discards a great deal of information that is likely to be useful for constructing dynamic user and word embeddings. Secondly, there are subgroups within the larger communities who are more and less interested in particular phenomena or events, such as the MeToo movement. Using individual user representations, rather than treating the movement as monolithic, allows for the emergence of subgroups with specialised interests. Finally, modelling users and words in the same space provides a strong platform for the analysis of extremist (and other) communities. In the next section, we provide a final exploration of this concept.
@@ 8 Splintering
As discussed in Section 2, the manosphere shows signs of macro-level splintering, with well-defined subgroups such as Incels forming more violent splinters. Here, we explore lower level splintering by investigating the clusters produced by our system. For a clearer visualisation, we use K-means clustering with $k=50$. To estimate the level of violent language per cluster, we use a subset of 39 terms from the Incel Violent Extremism Dictionary of Baele et al. [4] that specifically references physical violence (e.g. kill, rape, murder). We find the dot product of each cluster with the centroid of the violence lexicon at every timestep.
The results are shown in Figure 4. Each vertically-grouped collection represents the violence scores for a specific cluster at different timesteps. As per Section 7, the cluster embeddings are fixed, such that variations are caused by changes in word embeddings over time. The marker size corresponds to the cluster size. For reference, the community at index 15 consists of 48,993 users.
For the manosphere communities, we observe a trend of larger, less violent groups and smaller, more violent groups within each category. This supports the existence of low-level splintering; i.e. subgroups that are more violent than the mainstream within each category.
The Incel clusters have the highest mean violence score, which supports the idea that they are a more extreme splinter of the manosphere. The MRAs also show a penchant for violent language, however, this may be due to their interest in discussing sexual assault accusations. The cluster associated with Pick-up Artists has relatively low scores, which reflects their main interest being seduction tricks.
As discussed in Section 6.1, the MGTOW community yields the most subclusters. Future work may investigate this result, as it could be indicative of a more fractured subcommunity structure. We note that the largest MGTOW cluster has a very low violence score, which could be related to the notion that these communities tend to be very supportive and empathetic of their own members [21].
We observe that the mental health clusters have relatively low violence scores. One mental health community has a slightly higher than average score. Upon investigation, we note that this cluster has substantial overlap with the r/suicidewatch community.
@@ 9 Conclusion
In this work, we introduce a novel architecture for modelling online communities. Experimental results show that the resulting embeddings yield better results in clustering and embedding forecasting evaluations. Beyond the quantitative improvement, our analyses in Sections 7 and 8 illustrate the usefulness of the learned embeddings in the context of an online extremist community. These analyses enable novel insights into the manosphere and its subgroups, and have the potential to support monitoring of these groups. The question of how to act upon such information is less clear; past suppression attempts have resulted in splinter groups migrating to specialised platforms. Continued cross-disciplinary efforts remain essential in mitigating online harms.
@@ Appendix A Preliminary experiments
Content reconstruction error Adjacency reconstruction error
Model >0-MAE 0-MAE Weighted MAE >0-MAE 0-MAE WMAE
Chimera 2.650 2.551 2.675 2.549 2.543 2.574
- nonneg. constraint 2.607 1.573 2.623 1.566 1.566 1.582
+ scaling zeros 2.365 1.742 2.383 1.723 1.591 1.739
+ missing user masking 2.332 2.113 2.353 1.703 1.693 1.720
+ biases 2.350 2.225 2.372 1.693 1.795 1.711
Table 5: Results for the preliminary experiments.
@@ Appendix B Community prediction model
The community prediction model comprises one LSTM layer with 256 units, followed by two dense layers of 512 and 256 units, respectively. We use a grid search to determine the best dropout and learning rate values for each model, experimenting with $\alpha\in\{0.001,0.01,0.1\}$ and $p\in\{0.1,0.2,0.5\}$. The mean squared error is used as training loss, and Adam [18] is used for optimisation.

### v2

@@ 4.3 Data limitations
5 Evaluation
In our evaluation, we aim to isolate the effect of modelling time, social dynamics and language, compared to models that neglected some of these interactions. We therefore follow and expand the evaluation setting of Appel et al. (2019), using two test cases for evaluation: clustering and forecasting. For the latter, we perform an embedding prediction task as well as predicting with which communities a user will interact.
@@ 5.1 Clustering evaluation
K-means clustering based on the temporal user embeddings $U$ is used to discover sub-communities. Following Appel et al. (2019), these clusters are calculated across different timesteps, meaning that a user may belong to the same cluster or to different clusters over different timesteps. Intuitively, the clusters may be thought of as topics, such that users who do not overlap temporally but have similar embeddings at different timesteps may be assigned to the same cluster.
Using the subreddit structure, we evaluate the soundness of these clusters based on cluster purity, a widely-used metric for evaluating uniformity of cluster elements as the fraction of users in each cluster that belong to the dominant class (Schütze, Manning, and Raghavan 2008). Since one user may be associated with multiple subreddits at a given timestep, we use a multilabel interpretation of cluster purity, meaning that the dominant class is the class that has the highest representation in a cluster. Let $L(i,t)$ denote the label set of user $i$ at time $t$, $\Omega=\{\omega_{1},\omega_{2},\ldots,\omega_{K}\}$ the set of clusters, and =$\mathbb{C}=\{c_{1},c_{2},\ldots,c_{J}\}$ the set of classes. We report the average cluster purity over $K$ clusters, given:
@@
The class labels comprise a two-level hierarchy, consisting of a more general category label (with 7 classes: Incels, MGTOW, TRP, PuA, MRA, mental health and criticism) and a more finegrained subreddit label (with 50 classes). Allowing for some subgroup specialisation at different levels of granularity, we report results for $k\in\{10,100,1000\}$, using both subreddit and category-level labels. Since the clustering algorithm is dependent on random initialisation, we report the mean and standard deviation over 5 runs.
K=10 K=100 K=1000
Model Category Subreddit Category Subreddit Category Subreddit
@@ 5.3 Community prediction
We use a distance-based approach for this purpose. We find the centroid of each community, i.e. the mean of the embeddings of all users who interacted with said community in $t_{1}$ to $t_{T}$. We then calculate the cosine similarity $s$ between the predicted user embedding and the centroid of each community. The inferred user label is given by the max-normalised similarities to all centroids, i.e. $\hat{y_{i,t}}=\frac{s_{i,t}}{max(s_{i,t})}$, to capture the relative relevance of each community to a user.
@@
This formulation equates to a multilabel regression task, which we evaluate using the concordance index (CI; Harrell et al. 1982). The CI is the fraction of concordant predictions (i.e.  ($y_{i}>y_{j}$, $\hat{y}_{i}>\hat{y}_{j}$) or ($y_{i}<y_{j}$, $\hat{y}_{i}<\hat{y}_{j}$); in other words, a pairwise comparison is performed on all possible pair permutations, and the fraction that are correctly ordered is calculated. We report both the mean per-class CI (comparing predictions across users for a specific community) and the within-sample CI (evaluating the orderings per user).
6 Implementation details
We extend the TensorFlow implementation of Appel et al. (2019) to implement the proposed model. To perform hyperparameter tuning and to establish the efficacy of the four training optimisations outlined in Section 3.4, we run preliminary experiments on a reduced training set. These results, reported in Appendix A, show that the optimisations provide statistically significant improvements in the reconstruction accuracy in a smaller scale evaluation setting, and as such they are incorporated in all models that are compared in this work. Our code and models will be released to the community.
The asymptotic complexity of our method matches that of the approach proposed by Appel et al. (2019). The introduction of temporal word embeddings in our method adds ($T-1$) additional matrix multiplications to the objective function, where $T$ is a constant scalar. Since the number of additional operations is proportional to $T$, the overall asymptotic complexity of our method remains unchanged. The Cerberus system was trained on a Google Cloud instance with 256GB memory and 64CPUs for approximately 1 week. Reducing the training costs would be trivial if required, as the interaction matrices are sparse and do not need to be read into memory in full. The vocabulary could also be reduced by excluding uncommon words or stopwords. Additionally, GPUs can be explored for faster training. In these evaluations, we keep the implementation as-is to replicate the experimental setting used in the precursor works.
@@ 7.1 Cluster purity
Cluster purity over $K$ clusters is shown in Table 2. Our model outperforms the StatCont and NoAdj systems in 4 out of 6 cases, with the largest improvement observed for subreddit purity at $K=1000$. The NoAdj model has the same score as ours for category purity at $K=100$, whereas the StatCont model performs the same as ours for category purity at $K=1000$. We note that cluster purity is better at larger $K$, which is to be expected; purity will be 1 if $K=N$, but the comparison of different models at a given $K$ gives us a useful comparison point. In all cases, the cluster purity is higher for the category labels than the subreddit labels, which again is to be expected; since users are distributed over more classes, clusters are less likely to be dominated by a single subreddit.
We also compare our cluster purity to embeddings that are obtained through standard static (i.e. time-aggregated) matrix factorisation of the content embedding (MatFact) and static shared matrix factorisation of content and adjacency (SharedMF). In this case, only one embedding is produced per user and therefore a user is assigned to only one cluster. We calculate purity based on all communities with which a user interacted over the training period. From the results, we note that the cluster purity is lower in this setting. A possible reason for this result is that users vary in their category associations over time, which is not captured well by the time-aggregated clustering. The static shared matrix factorisation model outperforms the static model without social information.
The per-cluster purity for $K=10$, at the category level and using Cerberus, is shown in Table 3, sorted by cluster size. Size is based on the number of (user, timestep) tuples in the cluster. We note that there is a substantial difference between the high-scoring and low-scoring clusters, with clusters 2, 7, 8, 9 and 10 being high-purity clusters representing 4 out of the 7 higher level categories in the data. With the exception of cluster 2, these are all relatively small clusters, possibly representing the dedicated core users of each ideology who do not engage with other categories. Cluster 2 is a large and high-purity cluster which represents the mental health category. This provides support for the soundness of the clustering and embeddings, as we would expect this control group to be disparate.
@@ 7.2 Evolution of users
Figure 2: Predicting user evolution over time. Grey markers represent the trajectory of a given user, whereas coloured dots are community centroids. At each timestep, the predicted next embedding is shown in red and the true next embedding in black, while the data seen by the model is in grey.
In the context of extremist groups, the ability to represent a user as a dynamic entity at different timesteps is a key benefit of the temporal factorisation approach. Prior work (De Kock and Hovy 2024; Ferrara et al. 2016) developed predictive models to anticipate whether a user will engage with an extremist group; however, these works relied on derived indicators, whereas we our system models users and communities in the same embedding space.
@@
As described in Section 5, we evaluate user evolution modelling in two ways: embedding prediction (comparing learned embeddings to predicted embeddings, within the training window) and community prediction (predicting beyond the training window and comparing to true interactions). These results are shown in Table 4. Though we use the subreddit structure to evaluate our models, they can be applied in forums without explicit subcommunities.
@@ 7.3 Evolution of words
Figure 3: Word relevance to different clusters over time.
@@
There are three main takeaways from these figures. Firstly, dynamic word embeddings are important: a static model would produce an aggregate representation of Kavanaugh, which discards a great deal of information that is likely to be useful for constructing informative user, word and community embeddings. Secondly, there are subgroups within the larger communities who are more and less interested in particular phenomena or events, such as the MeToo movement. This is captured by our model by using individual user representations, rather than treating the movement as monolithic, allowing for the emergence of subgroups with specialised interests. Finally, modelling users and words in the same space provides a strong platform for the analysis of extremist (and other) communities. In the next section, we provide a final exploration of this concept.
@@ 8 Splintering
As discussed in Section 4.1, the manosphere shows signs of macro-level splintering, with well-defined subgroups such as Incels forming more violent splinters. Here, we explore lower-level splintering by investigating the clusters produced by our system. For a clearer visualisation, we use K-means clustering with $k=50$.
To estimate the level of violent language per cluster, we use a subset of 39 terms from the Incel Violent Extremism Dictionary of Baele, Brace, and Ging (2023) that specifically references physical violence (e.g. kill, rape, murder) and find the centroid of their Cerberus embeddings. A similar approach for finding a concept centroid by averaging word embeddings is used by Mendelsohn, Tsvetkov, and Jurafsky (2020). We then find the dot product of each cluster with the violence centroid at every timestep. A limitation of this approach is that it does not consider the context in which a word is used (for example, sarcasm); however, by relying on embedding distances of a cluster vector to a concept vector, instead of word counts as used in prior work (Baele, Brace, and Ging 2023), we can obtain a more nuanced measure of the general tendency within the cluster to use language related to violence. While the centroid is constructed based on words that unambiguously relate to physical violence, it should also be close to the representations of emerging slang words that implicitly signal violence within the community.
The results are shown in Figure 4. Each vertically-grouped collection represents the violence scores for a specific cluster at different timesteps. As per Section 7.3, the cluster embeddings are fixed, such that variations are caused by changes in word embeddings over time. The marker size corresponds to the cluster size. For reference, the community at index 15 consists of 48,993 user representations.
For the manosphere communities, we observe a trend of larger, less violent groups and smaller, more violent groups within each category. This supports the existence of low-level splintering; i.e. subgroups that are more violent than the mainstream within each category. Future work may investigate the emergence and focuses of these smaller and more violent groups.
The Incel clusters have the highest mean violence score, which supports the idea that they are a more extreme splinter of the manosphere. The MRAs also show a penchant for violent language, however, this may be due to their interest in discussing sexual assault accusations. This highlights another potential area of future work: accounting for polysemy and linguistic context in the content representation. We observe that the mental health clusters generally have relatively low violence scores, with the exception of one ($i=5$) which has a slightly higher than average score. Upon investigation, we note that this cluster has substantial overlap with the r/suicidewatch community. The cluster associated with Pick-up Artists ($i=44$) also has relatively low scores, which reflects their main interest being seduction tricks.
As discussed in Section 7.1, the MGTOW community yields the most subclusters. Future work may investigate this result, as it could be indicative of a more fractured subcommunity structure. We note that the largest MGTOW cluster has a very low violence score, which may be related to the notion that these communities tend to be very supportive and empathetic towards their own members (Rich and Bujalka 2023).
@@ 9 Conclusion
In this work, we introduce a novel architecture for modelling online communities. Experimental results show that the resulting embeddings yield better results in clustering and embedding forecasting evaluations. Beyond the quantitative improvement, our analyses in Sections 7.3 and 8 qualitatively illustrate the usefulness of the learned embeddings in the context of an online extremist community. These analyses enable novel insights into the manosphere and its subgroups, and have the potential to support monitoring of these groups. The question of how to act upon such information is less clear and should be treated with caution; past suppression attempts have resulted in splinter groups migrating to specialised platforms. False accusations may have a devastating impact on the lives of those impacted; as such, our approach does not attempt to classify users as “radicalised” or “extremist” as previous works do, but instead utilises relative distances to communities. Continued cross-disciplinary efforts remain essential in mitigating online harms.
@@ Appendix A Preliminary experiments
Content reconstruction error
Model >0-MAE 0-MAE WMAE
Chimera 2.650 2.551 2.675
- nonneg. constraint 2.607 1.573 2.623
+ scaling zeroes 2.365 1.742 2.383
+ missing user masking 2.332 2.113 2.353
+ biases 2.350 2.225 2.372
Adjacency reconstruction error
Model >0-MAE 0-MAE WMAE
Chimera 2.549 2.543 2.574
- nonneg. constraint 1.566 1.566 1.582
+ scaling zeroes 1.723 1.591 1.739
+ missing user masking 1.703 1.693 1.720
+ biases 1.693 1.795 1.711
Table 5: Results for the preliminary experiments.
@@ Appendix B Community prediction model
The community prediction model comprises one LSTM layer with 256 units, followed by two dense layers of 512 and 256 units, respectively. We use a grid search to determine the best dropout and learning rate values for each model, experimenting with $\eta\in\{0.001,0.01,0.1\}$ and $p\in\{0.1,0.2,0.5\}$. The mean squared error is used as training loss, and Adam (Kingma and Ba 2014) is used for optimisation.